that means of the functions f /g where f x, y is equal to 2 x y and g x, y is equal to x square + y square is r square which statement is true or false give reason
1
Expert's answer
2020-06-21T15:04:51-0400

"f(x,y) = 2xy, g(x,y) = x^2 + y^2"


for "\frac{f(x,y)}{g(x,y)}" "= \frac {2xy}{(x^2 + y^2)}"


This function is defined at all values of x and y except "x=y=0" .


So its domain cannot be "\R^2" .